Q : 1 The distance x meters moved by a partide int seconds is given by x = t + 3t + 4 . Find the velocity and accelerating after 3 seconds .

Q : 2 The radius of a circle is increasing uniformly at the rate of 3 cm per second . Find the rate at which its area is increasing when radius is 10cm .

Q : 3 If displacement is s = sin2t , find , its acceleration .


1
Expert's answer
2021-05-24T15:27:37-0400

Q : 1 ) x = t + 3t + 4 ........................Equation(1)


We know that velocity and acceleration are given by:


v = "\dfrac{dx}{dt}"


a = "\dfrac{d^2x}{dt^2}"


Differentiating equation (1) with respect to t, we have


v = "\dfrac{dx}{dt}" = 4 ............................................Equation (2)



Differentiating Equation (2) with respect to t, we have


a = "\dfrac{d^2x}{dt^2}" = 0



Hence, we see that the velocity at any time is constant and is equal to 4 m/s and acceleration is equal to 0.



Q : 2) "\dfrac{dr}{dt}" = 3 cm/s


Now, area of a circle is given by

A = "\pi" "*" r2 ..................Equation(3)


Differentiating the above equation with respect to t, we have


"\dfrac{dA}{dt}" = "\pi" "*" 2 "*" r "*" "\dfrac{dr}{dt}"



Substituting the value of r and "\dfrac{dr}{dt}" in the above equation, we have



"\dfrac{dA}{dt}" = "\pi" "*" 2 "*" 10 "*" 3 cm2 / s



"\dfrac{dA}{dt}" = 60"\pi" cm2/s = 188.4 cm2 / s




Q : 3) s = "\sin2t" ................Equation(4)


Now, acceleration is given by


a = "\dfrac{d^2s}{dt^2}"



Differentiating Equation (1) twice with respect to t, we have


a = "\dfrac{d^2s}{dt^2}" = "\dfrac{d}{dt}(2cos 2t)"


a = "\dfrac{d^2s}{dt^2}" = - 4 "\sin2t"


a = - 4 "\sin2t"